Questions 31-32: Consider the following table of data. We want to know if there is an association between gender and whether a college student is in a fraternity / sorority.
   Yes    No     Total
Male    12   39   51
Female    33   76   109
Total    45   115   160


31. Find the expected number of females that are not in a sorority (No).: *
(A) 30.6563
(B) 36.6563
(C) 76
(D) 78.3438

32. What are the degrees of freedom for this test?: *
(A) 1
(B) 3
(C) 44
(D) 159
Answer #1

expected number of females that are not in a sorority =

expected number E = = 78.3438

Hence option D is correct

32. Degree Of freedom DF :

DF = (r -1) *(c-1)

DF = (2-1)*(2-1)

DF = 1*1 = 1

DF = 1

Hence Option A is correct